OpenStudy (darkbluechocobo):

Help with word problem Use functions to model and solve the following problem. The length of a rectangular football field is 14 meters more than twice the width. If the perimeter is 346 meters find the field's dimensions.

OpenStudy (anonymous):

length=l width=w perimeter=p so, l=2w+14 (2 times the width add 14) p=346 Perimeter will equal (2xwidth) +(2xlength) so 346=(2w)+(2l) since l=2w+14, 346=2w+2(2w+14) 346=2w+4w+28=6w+28 6w+28=346 6w=318 w=53, the width is 53 meters the length is 2w+14 l=(2x53)+14=120 the length is 120 meters.

OpenStudy (darkbluechocobo):

Oh thank you :p